# Nominal Evaluation of Automatic Multi-Sections Control Potential in Comparison To A Simpler One- or Two-Sections Alternative With Predictive Spray Switching

### v1

@@ 3.1 Comparison of 6 Experimental Setups
To address Problem 1 the two area coverage path planning strategies from Sect. 2.3, $\textsf{M}_{1}$ and $\textsf{M}_{2}$, are compared with 3 different sections control setups displayed in Fig. 2. Thus, overall 6 different setups are compared. The corresponding spray volumes are denoted $S_{\textsf{M}_{k}}^{j}$ with $k\in\{1,2\}$ and $j\in\{1,2,48\}$. The area coverage pathlengths for $\textsf{M}_{1}$ and $\textsf{M}_{2}$ shall be denoted by $L_{\textsf{M}_{1}}$ and $L_{\textsf{M}_{2}}$. Note that the pathlengths are the same for all 3 sections setups.
@@
Fig. 17: Visualisation of input data for 10 fields: coordinates of field contours and of obstacle areas if applicable. The first 5 fields are ordered in the upper row from left to right. The remaining 5 fields are ordered in the second row from left to right.
@@
The 10 real-world fields used for evaluation are depicted in Fig. 17. This data includes examples with multiple obstacle areas, freeform mainfield lanes, and non-convex field contours.
Quantitative results are summarised in Tables 2 and 3. Qualitative results for 2 examples are visualised in Fig. 18 and Fig. 19. Visualisations for all other examples are provided in A.
Ex. $A_{\text{field}}$ $L_{\textsf{M}_{1}}$ $L_{\textsf{M}_{2}}$ $\Delta L$ [m] $\Delta L$ $\%$]
@@
Table 3: Spray volume results for the 6 different setups $S_{\textsf{M}_{k}}^{j}$ with $k\in{1,2}$ and $j\in\{1,2,48\}$. For a given field area $A_{\text{field}}$ the theoretical ideal spray volume for uniform spraying is $S_{\text{field}}^{\text{ref}}=A_{\text{field}}s_{\text{volume}}^{\text{ref}}$. The differences measured in metres and percent are $\Delta S_{\textsf{M}_{k},\text{m}}^{j,\text{ref}}=S_{\textsf{M}_{k},\text{m}}^{j,\text{ref}}-S_{\text{field}}^{\text{ref}}$ and $\Delta S_{\textsf{M}_{k},\%}^{j,\text{ref}}=(S_{\textsf{M}_{k},\text{m}}^{j,\text{ref}}-S_{\text{field}}^{\text{ref}})/S_{\text{field}}^{\text{ref}}$, respectively.
@@ 4.1 Key Findings
    The pathlength savings potential for the optimisation-based $\textsf{M}_{2}$-method in comparison to the Boustrophedon-based $\textsf{M}_{1}$-method is between -3.9% and -12.3%, see Table 2. The optimisation-based $\textsf{M}_{2}$-method yielded better results throughout all experiments.
@@
    Both the 1- and 2-sections setup can be implemented robustly, and crucially and in contrast to the ASC-setup with 48 sections, without the need for any localisation sensors. This can be achieved by planting crops in a different direction along the headland and mainfield areas, respectively. Then, the intersection line between those 2 areas can serve as visual reference line at which spray switching commands can be triggered. In particular, this approach is also appropriate for manual operation. See Fig. 7 for illustration.
@@
In view of these findings plus the listing of the plethora of complexities associated with the spraying process in Table 1, the 1- and 2-sections setup present a reasonable alternative to ASC and offer multiple benefits. A recommended summary would be to (i) operate localisation-sensor free (no expensive RTK-GPS needed), (ii) use the $\textsf{M}_{2}$-method for area coverage path planning, (iii) prepare the intersection-line during the seeding process by planting crops in a different direction along the headland and mainfield areas, before (iv) switching manaually with one or two sections according to the methods $S_{\textsf{M}_{2}}^{1}$ or $S_{\textsf{M}_{2}}^{2}$.
@@ 4.2 Limitations
Limitations of the method are discussed. As stressed throughout the paper nominal conditions are assumed to evaluate the best-case performance that ASC can deliver. Thus, performance in any real-world conditions are degraded according to the effects listed in Table 1. However, the methods $S_{\textsf{M}_{2}}^{1}$ or $S_{\textsf{M}_{2}}^{2}$ are naturally also affected by uncertainties and inaccuracies in practice. For example, human driving experience is required to compensate for delays when switching at the intersection line. Nevertheless, the decisive advantages of the recommended method are the potential for (i) a sensor-free mechanical implementation, and (ii) robustness of the method.
@@ 5 Conclusion
This paper presented a nominal evaluation of Automatic Multi-Sections Control in comparison to a proposed simpler one- or two-sections alternative with predictive spray switching.
Four different hierarchical planning levels for area coverage were discussed. These include a path planning-, a block switching-, a sections switching- and a individual nozzle control-level. The focus of this paper was on the first three levels. Two area coverage path planning patterns, the Boustrophedon-pattern and an alternative pattern, were discussed. Optimised block switching-logics for each of these two patterns were presented. Three different variations on the sections-switching level were compared.
@@
In contrast to ASC, a preferred method is suggested that (i) minimises area coverage path length by employing a path planning technique different from Boustrophedon-based path planning, (ii) offers intermediate overlap, (iii) crucially and in contrast to ASC can be implemented sensor-free (low-cost) by visual inspection of the intersection line resulting when employing different planting directions in the mainfield and headland area, and (iv) is suitable for manual driving by following a pre-planned predictive spray switching logic and area coverage path plan.
The preferred method minimises pathlength, however requires overall more spray volume than a more complex ASC-method. An econominc preference between the methods ultimately boils down to the question if the benefits of pathlength savings (which imply Diesel and time savings) and cheaper spraying machinery (enabled by sensor-free operation) outweight higher input material costs caused by higher spray volumes.
For input material costs calculation it was highlighted that spray mixtures consist to a very large degree out of water and only to a small degree, often 0.5-2%, out of chemical component (herbicides, pesticides and the like). Assuming typically much cheaper costs for water than for the chemicals, aforementioned increase in higher spray volumes does not imply the same percentage-increase in input material costs but is typically much less.

### v2

@@ 3.1 Comparison of 6 Experimental Setups
To address Problem 1 the two area coverage path planning strategies from Sec. 2.3, $\textsf{M}_{1}$ and $\textsf{M}_{2}$, are compared with 3 different sections control setups in Fig. 2. Thus, overall 6 different setups are compared. The corresponding spray volumes are denoted $S_{\textsf{M}_{k}}^{j},~\forall k\in\{1,2\},~\forall j\in\{1,2,48\}$. The area coverage pathlengths for $\textsf{M}_{1}$ and $\textsf{M}_{2}$ shall be denoted by $L_{\textsf{M}_{1}}$ and $L_{\textsf{M}_{2}}$. Note that the pathlengths are the same for all 3 sections setups.
@@
Fig. 17: Visualisation of input data for 10 fields: coordinates of field contours and obstacle areas if applicable. The first 5 fields are ordered in the upper row from left to right. The remaining 5 fields are ordered in the second row from left to right.
@@
The 10 real-world fields used for evaluation are depicted in Fig. 17. This data includes examples with multiple obstacle areas, freeform mainfield lanes, and non-convex field contours. Quantitative results are given in Table 2 and 3. Qualitative results for 2 examples are visualised in Fig. 18 and 19. Visualisations for all other examples are provided in A.
Ex. $A_{\text{field}}$ $L_{\textsf{M}_{1}}$ $L_{\textsf{M}_{2}}$ $\Delta L$ (m) $\Delta L$ ($\%$)
@@
Table 3: Spray volume results for 6 different setups, $S_{\textsf{M}_{k}}^{j},~\forall k\in\{1,2\},~\forall j\in\{1,2,48\}$. For a given field area $A_{\text{field}}$, the theoretical ideal spray volume for uniform spraying is $S_{\text{field}}^{\text{ref}}=A_{\text{field}}s_{\text{volume}}^{\text{ref}}$. The differences measured in metres and percent are $\Delta S_{\textsf{M}_{k},\text{m}}^{j,\text{ref}}=S_{\textsf{M}_{k},\text{m}}^{j,\text{ref}}-S_{\text{field}}^{\text{ref}}$ and $\Delta S_{\textsf{M}_{k},\%}^{j,\text{ref}}=(S_{\textsf{M}_{k},\text{m}}^{j,\text{ref}}-S_{\text{field}}^{\text{ref}})/S_{\text{field}}^{\text{ref}}$, respectively.
@@ 4.1 Key Findings
    The pathlength savings potential for the optimisation-based $\textsf{M}_{2}$-method in comparison to the Boustrophedon-based $\textsf{M}_{1}$-method is between -3.9% and -12.3% (see Table 2). $\textsf{M}_{2}$ yielded better results throughout all experiments. Since path planning precedes spraying according to the hierarchical structure of Fig. 6, it can be designed independently from the spraying technique.
@@
    Both the 1- and 2-sections setup can be implemented robustly, and crucially and in contrast to the ASC-setup, without the need for any localisation sensors. This can be achieved by planting crops in a different direction along the headland and mainfield areas, before using the intersection line between those 2 areas as visual reference line at which spray switching commands can be triggered. This approach is also appropriate for manual operation. See Fig. 7 for illustration.
@@
In view of these findings plus the listing of the plethora of complexities associated with the spraying process in Table 1, the 1- and 2-sections setup present a reasonable alternative to ASC and offer multiple benefits. A recommended summary would be to (i) operate localisation-sensor free (no expensive RTK-GPS needed), (ii) use the $\textsf{M}_{2}$-method for area coverage path planning, (iii) prepare the intersection-line during the seeding process by planting crops in a different direction along the headland and mainfield areas, before (iv) switching manually with one or two sections according to the methods $S_{\textsf{M}_{2}}^{1}$ or $S_{\textsf{M}_{2}}^{2}$. An economic evaluation is provided next.
4.2 Economic cost discusstion
Let the spray volume difference between a 1- or 2-sections solution and a multi-sections solution (ASC) be denoted by
$\Delta S_{\textsf{M}_{k}}^{j,\text{ASC}}=S_{\textsf{M}_{k}}^{j}-S_{\textsf{M}_{k}}^{48},~\forall k\in\{1,2\},~\forall j\in\{1,2\}.$ (6)
Then, assuming a water ratio of 99% for the spray mixture, the corresponding cost difference is
$\Delta C_{\textsf{M}_{k}}^{j,\text{ASC}}=(0.01\cdot C_{\text{chemical}}+0.99\cdot C_{\text{water}})\Delta S_{\textsf{M}_{k}}^{j,\text{ASC}},$ (7)
$\forall k\in\{1,2\},~\forall j\in\{1,2\}$.
Let the hectares-normalised values for (6) and (7) be denoted by $\Delta\bar{S}_{\textsf{M}_{k}}^{j,\text{ASC}}$ and $\Delta\bar{C}_{\textsf{M}_{k}}^{j,\text{ASC}}$ and measured in ($\ell$/ha) and (EUR/ha), respectively. Let the purchase cost difference in EUR between a ASC-machinery and a simpler 1- or 2-sections alternative be denoted as $\Delta K_{\text{ASC}}$. Then, two economic cost discussions are given.
First, ASC promises to save spray volume. ASC is profitable in comparison to a simpler 1- or 2-sections solution if
$\Delta K_{\text{ASC}}\leq\Delta\bar{C}_{\textsf{M}_{k}}^{j,\text{ASC}}.$ (8)
[…]
Second, an alternative to evaluate the economic potential of ASC can be provided by the following second inequality,
$\Delta K_{\text{ASC}}\leq\Delta\bar{C}_{\textsf{M}_{k}}^{j,\text{ASC}}A_{\text{total}}N_{\text{runs}}^{\text{field}}N_{\text{years}},$ (9)
where $A_{\text{total}}$ denotes total farm-wide field area in hectares, $N_{\text{runs}}^{\text{field}}$ the number of field runs per year required for spraying, and $N_{\text{years}}$ a number of years, respectively. The reformulation of (9) permits to calculate after how many years the spray volume savings of an ASC-solution economically outweigh a simpler 1- or 2-sections alternative. The number of years at which equality is attained in (9) shall be denoted by $N_{\text{years}}^{\star}$.
To evaluate (9) multiple parameters have to be chosen. $\Delta\bar{C}_{\textsf{M}_{k}}^{j,\text{ASC}}$ is computed based on data, by (i) computing the average $\Delta\bar{S}_{\textsf{M}_{k}}^{j,\text{ASC}}$ from the 10 field examples according to Table 4 before (ii) computing its corresponding cost difference $\Delta\bar{C}_{\textsf{M}_{k}}^{j,\text{ASC}}$ according to (7). Values for $C_{\text{water}}$, $C_{\text{chemical}}$, $A_{\text{total}}$ and $N_{\text{runs}}^{\text{field}}$ are selected according to Table 5. Results for the evaluation of (9) are presented in Table 6.
[…]
Second, for larger farm sizes of $A_{\text{total}}=1000$ha the situation changes. For example, here it is $N_{\text{years}}^{\star,30}=2.2$ years.
Overall, Table 6 demonstrates that different values for $A_{\text{total}}$, $C_{\text{chemical}}$ and $\Delta K_{\text{ASC}}$ largely shape after how many years the spray volume savings of an ASC-solution economically outweigh a simpler 1- or 2-sections alternative. The smaller the farm, the smaller the cost for chemicals and the larger $\Delta K_{\text{ASC}}$, the less ASC is economically worth it.
Ex. $A_{\text{field}}$ $\Delta S_{\textsf{M}_{1},\ell/\text{ha}}^{1,\text{ASC}}$ $\Delta S_{\textsf{M}_{1},\ell/\text{ha}}^{2,\text{ASC}}$ $\Delta S_{\textsf{M}_{2},\ell/\text{ha}}^{1,\text{ASC}}$ $\Delta S_{\textsf{M}_{2},\ell/\text{ha}}^{2,\text{ASC}}$
(ha) ($\ell$/ha) ($\ell$/ha) ($\ell$/ha) ($\ell$/ha)
1 6.0 11.7 8.9 13.8 10.5
2 14.6 12.7 10.5 14.0 10.9
3 10.1 18.7 15.8 19.0 13.5
4 11.8 12.0 8.1 13.5 8.8
5 13.5 10.5 8.9 8.6 6.9
6 38.5 19.8 19.0 19.9 19.0
7 6.5 26.9 26.5 27.8 26.9
8 13.6 22.9 19.8 30.9 22.6
9 4.5 38.4 32.6 56.3 50.0
10 7.2 12.7 9.4 12.9 9.7
Avg. 12.6 18.6 16.7 22.5 18.7
Table 4: The hectares-normalised spray volumes that a 1- or 2-sections solution requires more than ASC. Data is deduced from Table 3.
Parameter Value Unit
[…]
$A_{\text{total}}$ $\in\{30,100,300,600,1000\}$ ha
$N_{\text{runs}}^{\text{field}}$ 8 -
Table 5: Parameters used for economic cost calculations.
[…]
$\Delta K_{\text{ASC}}=100000$EUR
$A_{\text{total}}$ $N_{\text{years}}^{\star,30}$ $N_{\text{years}}^{\star,10}$
30 74.2 219.7
100 22.3 65.9
300 7.4 22.0
600 3.7 11.0
1000 2.2 6.6
$\Delta K_{\text{ASC}}=200000$EUR
$A_{\text{total}}$ $N_{\text{years}}^{\star,30}$ $N_{\text{years}}^{\star,10}$
30 148.4 439.3
100 44.5 131.8
300 14.8 43.9
600 7.4 22.0
1000 4.5 13.2
[…]
$\Delta K_{\text{ASC}}=100000$EUR
$A_{\text{total}}$ $N_{\text{years}}^{\star,30}$ $N_{\text{years}}^{\star,10}$
30 82.6 244.7
100 24.8 73.4
300 8.3 24.5
600 4.1 12.2
1000 2.5 7.3
$\Delta K_{\text{ASC}}=200000$EUR
$A_{\text{total}}$ $N_{\text{years}}^{\star,30}$ $N_{\text{years}}^{\star,10}$
30 165.2 489.3
100 49.6 146.8
300 16.5 48.9
600 8.3 24.5
1000 5.0 14.7
[…]
$\Delta K_{\text{ASC}}=100000$EUR
$A_{\text{total}}$ $N_{\text{years}}^{\star,30}$ $N_{\text{years}}^{\star,10}$
30 61.3 181.6
100 18.4 54.5
300 6.1 18.2
600 3.1 9.1
1000 1.8 5.4
$\Delta K_{\text{ASC}}=200000$EUR
$A_{\text{total}}$ $N_{\text{years}}^{\star,30}$ $N_{\text{years}}^{\star,10}$
30 122.6 363.2
100 36.8 109.0
300 12.3 36.3
600 6.1 18.2
1000 3.7 10.9
[…]
$\Delta K_{\text{ASC}}=100000$EUR
$A_{\text{total}}$ $N_{\text{years}}^{\star,30}$ $N_{\text{years}}^{\star,10}$
30 73.8 218.5
100 22.1 65.5
300 7.4 21.8
600 3.7 10.9
1000 2.2 6.5
$\Delta K_{\text{ASC}}=200000$EUR
$A_{\text{total}}$ $N_{\text{years}}^{\star,30}$ $N_{\text{years}}^{\star,10}$
30 147.6 437.0
100 44.3 131.1
300 14.8 43.7
600 7.4 21.8
1000 4.4 13.1
[…]
Above results are not flattering with respect to the economic profitability of ASC. Therefore, the natural question arises: How inaccurate is above parameter-based economic cost discussion?
The cost of pesticides (including herbicides, fungicides, and insecticides) for cereal crop production can vary widely based on several factors, including the type of persticide used, regional market conditions, the size of the farm, and the specific application needs (e.g., pre-emergence, post-emergence products). It’s also important to note that pesticide costs can vary greatly depending on the region, with some countries offering subsidies or differing regulations that might affect price points. Additionally, larger-scale operations might benefit from bulk discounts. Thus, overall parameter $C_{\text{chemical}}$ is difficult to estimate since it can regionally greatly vary. Above provided two inequalities, (8) and (9), permit a reader and practitioner to quickly evaluate their own cost models for $C_{\text{chemical}}$, and similarly for $C_{\text{water}}$. In contrast to $C_{\text{chemical}}$, $A_{\text{total}}$ and $N_{\text{runs}}^{\text{field}}$ are well measurable.
The water ratio is very influential on results. A water ratio of 99% was assumed for the spray mixture (see (7)). If this changed to 98% then $N_{\text{years}}^{\star}$-results in Table 6 are approximately halved. Nevertheless, for $A_{\text{total}}=30$ha $N_{\text{years}}^{\star,30}$ would still in all scenarios in the order of multiple decades. Inversely, for a ratio larger than 99%, $N_{\text{years}}^{\star}$ would be increased.
$\Delta\bar{S}_{\textsf{M}_{k},\ell/\text{ha}}^{j,\text{ASC}}$ is data-based. It depends on the methods from Sec. 2.2-2.7 and field contour data. Parameters include $v_{\text{ref}}$, $W$, $w$ and $s_{\text{volume}}^{\text{ref}}$. Several comments can be made. First, traveling velocities and path are the same for ASC and the proposed 1- and 2-sections alternatives. Therefore nominally, $v_{\text{ref}}$ has no effect on above relative comparison of the methods. In practice, ASC becomes more difficult for higher velocities. To avoid coverage gaps (which are more important to avoid then spray overlaps) ASC has to be set more conservatively, which would result in overlaps and reduction in spray volume savings.
Second, the larger $w$ the more ASC converges to the 1- or 2-sections solution. The larger $W$ the less the number of mainfield lanes and number of turns overall. On the other hand, during turns higher velocities at the outer sections along the boom bar.
Third, a change in $s_{\text{volume}}^{\text{ref}}$ does not change switching signals but only the magnitude of spray volume. Thus, the coverage maps, e.g. in Fig. 18 and Fig. 19, do not change and thus also the relative difference between ASC and proposed 1- or 2-sections alternative does not change.
Fourth, above cost discussion is linear in parameters.
Finally, while the purchase price of an ASC-machinery is well measurable, the actual performance is in generally non-linear degraded by the effects listed in Table 1. Thus, above economic cost discussion is optimistic with respect to ASC since nominal optimal ASC was assumed for simulation experiments. This was done intentionally as outlined in Problem 1 to derive a best-case upper performance bound for ASC.
ASC-maintenance and software upgrade costs are not included in the cost analysis and would further degrade the economic potential of ASC, since these costs are expected to be higher than for a 1- or 2-sections sensor-free alternative.
To summarise, two inequalities (8) and (9) were used to evaluate the economic potential of ASC in comparison to a simpler 1- or 2-sections alternative. Overall, and assuming selected parameters according to Table 5 and 6, results strongly suggest that ASC is purely economically speaking (and this is decisive for many small farms) not profitable for at least small farms, and a simpler 1- or 2-sections alternative is recommended. Margins in Fig. 20 and Table 6 are simply too large to recommend otherwise. As long as costs $C_{\text{chemical}}$ are low and price difference $\Delta K_{\text{ASC}}$ high this statement will hold. Above provided two inequalities (8) and (9) permit a practitioner to quickly evaluate their own cost models for $C_{\text{chemical}}$ and $\Delta K_{\text{ASC}}$.
@@ 4.3 Limitations
Limitations of the method are discussed. As stressed throughout the paper, nominal conditions are assumed to evaluate the best-case performance that ASC can deliver. Thus, performance in any real-world conditions are degraded according to the effects listed in Table 1. However, the methods $S_{\textsf{M}_{2}}^{1}$ or $S_{\textsf{M}_{2}}^{2}$ are naturally also affected by uncertainties and inaccuracies in practice. For example, human driving experience is required to compensate for delays when switching at the intersection line. Nevertheless, the decisive advantages besides the economic benefits of the recommended method are the potential for (i) a sensor-free mechanical implementation, and (ii) robustness of this method.
@@ 5 Conclusion
This paper presented a nominal evaluation of Automatic Multi-Sections Control (ASC) in comparison to a proposed simpler one- or two-sections alternative with predictive spray switching. Four different hierarchical planning levels for area coverage were discussed. These include a path planning, a block switching, a sections switching and an individual nozzle control-level. The focus of this paper was on the first three levels. Two area coverage path planning patterns, the Boustrophedon-pattern and an alternative pattern, were discussed. Optimised block switching-logics for each of these two patterns were presented. Three different variations on the sections-switching level were compared.
@@
In contrast to ASC, a preferred method is suggested that (i) minimises area coverage path length by employing a technique different from Boustrophedon-based path planning, (ii) offers intermediate overlap, (iii) crucially and in contrast to ASC can be implemented sensor-free (low-cost) by visual inspection of the intersection line resulting when employing different planting directions in the mainfield and headland area, and (iv) is suitable for manual driving by following a pre-planned predictive spray switching logic and area coverage path plan.
The preferred method minimises pathlength, however requires overall more spray volume tha more complex and expensive ASC. A detailed economic cost analysis was provided. Fundamental to this discussion was that spray mixtures consist to a very large degree out of water and only to a small degree, often 0.5-2%, out of chemical component (herbicides, fungicides and the like), but water costs are typically much cheaper than the cost for chemicals. Surprisingly strong economic arguments were found to not recommend ASC for small farms.